OpenStudy (anonymous):

I have f(x)= x-2 square this is the symbolic form I have to make a table and graph it . How do I do this and what number do I start with

OpenStudy (radar):

Is it\[f(x)=(x-2)^{2}\] or \[f(x)=x-2^{2}\] I am going to assume it was the first rather than the latter.

OpenStudy (anonymous):

f(x)=\[\sqrt{x-2?}\]

OpenStudy (anonymous):

is X=-4

OpenStudy (radar):

(ok) I was wrong on both counts lol. For the graph i would use x at values of -2. -1, 0, 1, 2 x f(x) -2 2i Complex number -1 sq root of -3 again a complex value 0 sq root of -2 complex 1 sq root of -1 or i 2 0 3 1 4 sq root 2= 1.414 5 sq root 3= 1.73 6 2 that should allow you to start the graph.

OpenStudy (radar):

I don't think you can get a single solution with the info you have. Sorry but I got to run, will check later.

OpenStudy (anonymous):

THANK YOU

OpenStudy (radar):

You could say f(x) has real roots when\[x \ge2\]